OpenStudy (anonymous):

Prove that \[\int_0^\infty \frac{x\sin{x}}{1+\cos^2{x}}dx=\frac{\pi^2}{4}\]

OpenStudy (anonymous):

I don't know how to do this one. Do I need to use Riemman sums?

OpenStudy (anonymous):

Any idea guys?

OpenStudy (anonymous):

is that even true?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

Try it on wolframalpha and check.

OpenStudy (anonymous):

it seems that the indefinite integral is not elementary function

OpenStudy (anonymous):

This is bad =/.

OpenStudy (anonymous):

How do you know that @Tyifan12879?

OpenStudy (anonymous):

because i didnt work it out...

OpenStudy (binary3i):

integral of xf(x)dx with lim 0 to b = integral of b/2 f(x)dx. if the function f(x)=f(b-x)

OpenStudy (anonymous):

You mean \[\int_0^b xf(x)dx=\int_0^b\frac{b}{2}f(x)dx\]? @binary3i?

OpenStudy (anonymous):

if \[f(x)=f(b-x)\]

OpenStudy (binary3i):

yes but here b is \[\infty\]

OpenStudy (anonymous):

Yes that is a big problem =/

OpenStudy (anonymous):

Thinking...

OpenStudy (anonymous):

I give up for today. Thank you all!

OpenStudy (anonymous):

http://www.analyzemath.com/calculus/limits/squeezing.html You might be able to use this, though I'm not too sure

OpenStudy (turingtest):

Oh man, you think I can get into this school no-data? I'm already posting some of these questions in meta-math. Somebody tell me how to do these.

OpenStudy (turingtest):

mathmate's cooking up something good I hope...

OpenStudy (mathmate):

Let \[I\ =\ \int\limits\limits xsin(x)dx = \sin(x) - x \cos(x) +C\] which clearly diverges (and oscillates) as x approaches infinity. The denominator of the integrand, \[1\le (1+\cos^2(x)) \le 2\] makes the required integral behave similarly to I above. This leads me to believe that the given improper integral does not exist. Is it possible to check if there is a typo in either the integral or the result?

OpenStudy (turingtest):

I highly doubt it, this question is one of a series on an entrance exam to a rather prestigious school in Mexico. http://assets.openstudy.com/updates/attachments/4f03e1ade4b075b566519b63-no-data-1325655819441-cinvestaventranceexamjul2011.pdf I find it hard to imagine that there are mistakes like this on the test.

OpenStudy (turingtest):

Any ideas Zarkon?

OpenStudy (zarkon):

did you look at the limits of the question you posted?

OpenStudy (mathmate):

Did you notice that the upper limit is pi, and not infinity? So this integral at least exists!

OpenStudy (zarkon):

mathmate is correct with the problem in this thread

OpenStudy (turingtest):

I didn't post it, no-data did. That sure makes a lot more sense, but I still don't know how to go about it.

OpenStudy (zarkon):

I was letting you know that the above is not the same as the linked problem you posted

OpenStudy (mathmate):

The integral from 0 to pi checks with pi^2/4 using numerical integration. So back to the drawing boards, guys!

OpenStudy (turingtest):

Right, I just got the link myself so I'm not familiar with the problems. Should have checked though. I wonder how much time they give you for this test.

OpenStudy (zarkon):

is can be done using integration by parts

OpenStudy (turingtest):

as it is?

OpenStudy (zarkon):

yes...there is a little trick when evaluating the integral after parts is done once

OpenStudy (zarkon):

no...with pi...not infinity

OpenStudy (zarkon):

the integral does not converge when the upper limit is infinity

OpenStudy (turingtest):

no, I got that bit, mathmate made a convincing argument.

OpenStudy (zarkon):

use u=x and dv =the rest of the integrand

OpenStudy (turingtest):

ok thanks, I'll give it a try.

OpenStudy (turingtest):

Yeah, I don't know how to integrate\[\int\tan^{-1}(\cos x)dx\]is that even right so far?

OpenStudy (turingtest):

I mean I know I didn't include the other terms, but it seems to come to this integral

OpenStudy (zarkon):

you need the limits of integration

OpenStudy (turingtest):

but how can I use them if I have no idea how to integrate that?

OpenStudy (zarkon):

I said it is a trick :)

OpenStudy (zarkon):

you need them..they are very important

OpenStudy (anonymous):

I'm also stuck at \[\int \frac {x sin{x} dx}{1+cos^{2}{x}}=-x tan^{-1}{cos{x}}+\int tan^{-1}{cos{x}} dx\]

OpenStudy (turingtest):

I'm not going to sleep tonight :/

OpenStudy (zarkon):

you want a hint?

OpenStudy (turingtest):

I think I do

OpenStudy (anonymous):

So are you saying that to solve this I need to use the limits?

OpenStudy (zarkon):

do you know what this would be ... let f be an odd function what is \[\int\limits_{-a}^{a}f(x)dx\]

OpenStudy (turingtest):

0

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

you need to use that

OpenStudy (turingtest):

is arctan odd?

OpenStudy (zarkon):

it is

OpenStudy (turingtest):

no

OpenStudy (turingtest):

really?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

yes...arctan(cos(x)) is not...but you can fix that :)

OpenStudy (turingtest):

of course...

OpenStudy (turingtest):

Ok no more hints. This may not get solved by me tonight though.